OpenStudy (anonymous):

differentiate from the first principle x+1/x

OpenStudy (anonymous):

d/dx(x)=1 \[1/x=x ^{-1}\] then d/dx(x^-1)=-x^(x-2) so the final answer is: 1+-x^(x-2)

OpenStudy (anonymous):

Hey radar. Im your fan number 1

OpenStudy (radar):

Hey emunrradtvamg, thanks, hope you're having a nice day.